OpenStudy (anonymous):

I must use the equation A=Pe^rt. I am given the initial investment P=$7500 And I am given the amount of time that it will double (21years) I must find the annual %rate and the amount of money after ten years. Please use detailed step by step process so I can understand! (thanks so much!)

OpenStudy (anonymous):

\[2=e^{21r}\]solve for \(r\)

OpenStudy (anonymous):

in log form it is \[\ln(2)=21r\]so \[r=\frac{\ln(2)}{21}\]

OpenStudy (anonymous):

then a calculator http://www.wolframalpha.com/input/?i=log%282%29%2F21

OpenStudy (anonymous):

and then i guess turn your answer in to a percent by moving the decimal over two places left

OpenStudy (hlares):

Alright, so, we have P=7,500 and T=21 for the first part and we need to find R=annual percentage rate for the second part. Again, we will be using the natural logarithm to find and the fact that lnx=c and e^c=x. 15,000=(7,500)*e^(21r) 2=e^(21r) ln(2)=21r ln(2)/21=r r=0.03300... So, we will round the annual percentage rate to two significant decimal places: R=0.033 Now that we have R, we can find how much money we will have after ten years: A=(7,500)*e^(10*0.033) A=(7,500)*e^(0.33) A=10,432.261 Thus, we will have $10,432.26 after ten years.

OpenStudy (hlares):

I should add with the annual percentage rate: for the calculations, I left it as a decimal point. For the actual percent you will write down, multiply R=0.033 by 100 for R=3.3%.